OpenStudy (anonymous):

AP Calculus question: How to solve this without a calculator Sin(arctan(3/4))?

myininaya (myininaya):

\[\text{ Let } y=\arctan(\frac{3}{4}) => \tan(y)=\frac{3}{4} (=\frac{opposite}{adjacent} )\] Draw a right triangle.

myininaya (myininaya):

Find the hyp of that triangle and then find sin(y) and you are done.

OpenStudy (anonymous):

I got (3/5) hopefully that's the right answer! But thank you!

myininaya (myininaya):

That actually is right.